OpenStudy (anonymous):

Find the mass of 85mL sample of corn syrup, which has the density of 1.38 g/ml

OpenStudy (aaronq):

all you have to do is apply this formula. And by that, i mean plug in your values and solve algebraically: \(density=\dfrac{mass}{volume}\) can you identify in what category (mass, volume or density) the values they gave you belong to?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks for coming back on

OpenStudy (aaronq):

no problem, i do have to go soon. i got a test tmrw.

OpenStudy (anonymous):

well thnaks and good lucky on your test for tomorrow.. I also have a test tomorrow

OpenStudy (aaronq):

thanks ! good luck to you as well. so do you know where the values go?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i do

OpenStudy (aaronq):

so is the algebra part whats troubling you?

OpenStudy (anonymous):

yeah, but i kind figured it out by myself..

OpenStudy (anonymous):

it was just confusing

OpenStudy (aaronq):

yeah. just remember you have to perform the reverse operation. if you're dividing, multiply: \(x=\dfrac{y}{z}\rightarrow y=x*z\) if you're adding, subtract: \(x+y=z \rightarrow x=z-y\) good luck!

OpenStudy (anonymous):

thanks, i'll write this down..

OpenStudy (anonymous):

bye now GOOD LUCKY ON YOUR TEST

OpenStudy (aaronq):

thanks, see ya !

OpenStudy (anonymous):

bye